OpenStudy (anonymous):

Medal 4 help How many sides does a regular polygon have if one exterior angle is 1*? 120 180 360

OpenStudy (anonymous):

The complete perimeter comprises 360 degrees, therefore if exterior angle = 1°, number of sides = 360 [ Compare with a hexagon : 6 sides, therefore exterior angle = 360/6 = 60° or, alternatively, since exterior angle = 60°, then number of sides = 360/60 = 6 ]

OpenStudy (anonymous):

thnx:)

OpenStudy (anonymous):

welcome :)